1. Use a proportion to solve each problem.
1.1. the numerator of a certain fraction is 3 times the denominator.if the numerator is decreased by 1 and the denominator is increased by 2,the value of the resulting fraction is 5/2.find the original fraction.
1.2. thenjo can wash 6 chairs in 21 minutes.how many minutes would it take thenjo to wash 10 chairs?
1
Expert's answer
2020-10-08T14:31:46-0400

"(1)\\
\displaystyle\textsf{Let the fraction be}\hspace{0.1cm} \frac{x}{y}\\


\frac{x}{y} = \frac{3y}{y}\\
 

\displaystyle\frac{3y - 1}{y + 2} = \frac{5}{2}\\


\begin{aligned}
2(3y - 1) &= 5(y + 2)\\
6y - 2 &= 5y + 10\\
y &=10 + 2 = 12
\end{aligned}\\

\therefore \textsf{The original fraction is}\hspace{0.1cm} \frac{x}{y} = \frac{36}{12}.\\


(2)\\


6\hspace{0.1cm} \textsf{chairs can be washed in} \hspace{0.1cm}21\hspace{0.1cm} \textsf{minutes}. \\


1\hspace{0.1cm} \textsf{chair can be washed in} \hspace{0.1cm}\frac{21}{6}\hspace{0.1cm} \textsf{minutes}. \\

\therefore 10\hspace{0.1cm} \textsf{chairs can be washed in} \\\hspace{0.1cm}\frac{21\times 10}{6}= 35\hspace{0.1cm} \textsf{minutes}."